# Holographic entanglement negativity and thermodynamics in backreacted AdS black hole

### v1

@@ 5 HEN for adjacent subsystems
$x\equiv x^{1}=\biggl[-\frac{l_{j}}{2},\frac{l_{j}}{2}\biggr],\quad x^{i}=\biggl[-\frac{L}{2},\frac{L}{2}\biggr];\quad j=1,2,$ (5.1)
where i=2,3,,d1i=2,3,...,d-1 and $L\to\infty$. The holographic prescription then requires an appropriate combination of the entanglement entropies of the individual regions and their union [Jain:2017aqk, Jain:2017xsu]. This construction effectively isolates the quantum part of correlations by eliminating purely thermal contributions. In what follows, we will evaluate the holographic negativity for the adjacent strip configuration in the backreacted AdS black hole background and examine its dependence on quark density $\rho$, with particular focus on the low and high temperature limits, in order to contrast its behavior with that of the entanglement entropy discussed previously. Following the proposal given in [Jain:2017xsu], HEN for two adjacent subsystems is given by
$\mathcal{E}=\frac{3}{16G_{N}}(\mathcal{A}_{1}+\mathcal{A}_{2}-\mathcal{A}_{3}),$ (5.2)
Here, $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and $\mathcal{A}_{3}=\mathcal{A}_{1}+\mathcal{A}_{2}$ represent the areas of the corresponding RT surfaces, obtained using the prescription reviewed in section˜3. To express the entanglement negativity in terms of boundary parameters, one needs to determine the turning point $z_{t}$ as a function of the strip width $l$ by solving (3.6). However, as discussed in section˜3, this relation cannot be solved exactly for arbitrary $z_{t}$ and admits analytic solutions only in the low and high temperature regimes. Therefore, we evaluate the HEN in these two limiting cases using the corresponding results for the HEE.
@@ 5.1 Low temperature limit
The low temperature regime corresponds to the condition $z_{t}\ll z_{h}$. In terms of the boundary parameter, this limit translates to $Tl\ll 1$. By solving eq.˜3.6 perturbatively for $z_{t}$ in powers of $l$, one can obtain the corresponding expressions for the area functions $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and $\mathcal{A}_{3}$ in this regime, as outlined in [Chakrabortty:2020ptb]. The resulting expression for the area is given by
@@
$\begin{split}\mathcal{A}_{j}=\frac{2}{d-2}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}&+R^{d-1}\left(\frac{L}{l_{j}}\right)^{d-2}S_{0}\Biggl(1+\rho S_{2}\left(\frac{4\pi T_{f}l}{d}\right)^{d-1}\\ &+(1-\rho)S_{1}\left(\frac{4\pi T_{f}l_{j}}{d}\right)^{d}+\mathcal{O}\left((T_{f}l_{j})^{2d-2}\right)\Biggr).\end{split}$ (5.3)
@@
Here, the index $j=1,2,3$ labels the subsystems $A_{1}$, $A_{2}$, and $A_{3}$, corresponding to their respective strip widths. In eq.˜5.3, the first term with $\epsilon$ denotes the divergent contribution, while $S_{0}$, $S_{1}$, and $S_{2}$ are constants depending only on the spacetime dimension $d$ and are given in eq.˜3.10.
Using (5.2) together with the low-temperature expansion eq.˜5.3 for the area functionals, the HEN for two adjacent strip intervals in the deformed AdS black hole background takes the form
@@
$\begin{split}\mathcal{E}&=\frac{3}{8G_{N}(d-2)}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}\quad+\frac{3R^{d-1}L^{d-2}}{8G_{N}}S_{0}\Biggl(l_{1}^{2-d}+l_{2}^{2-d}-(l_{1}+l_{2})^{2-d}\\ &\qquad\qquad\qquad\qquad\qquad\qquad-\;2S_{1}(1-\rho)\,l_{1}l_{2}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{\!d}+\mathcal{O}\!\big((T_{f}l_{j})^{2d-2}\big)\Biggr).\end{split}$ (5.4)
From the above expression eq.˜5.4, it follows that :
@@
  • The first term is the standard ultraviolet divergence arising from short-distance correlations near the entangling surface. The leading finite contribution (independent of $T_{f}$) is the familiar vacuum-like short-distance scaling for adjacent strip geometries and encodes the purely geometric contribution to negativity.
@@
  • The first nontrivial thermal/backreaction correction appears at order $T_{f}^{\,d}$ and is multiplied by the factor $(1-\rho)$. In low temperature regime $T_{f}^{\,d}$ term is the leading thermal correction; higher-order terms scale as $(T_{f}l_{j})^{2d-2}$ and are subleading for sufficiently small $T_{f}l_{j}$.
  • The correction proportional to $(1-\rho)T_{f}^{\,d}$ reduces the finite part of $\mathcal{E}$. For fixed $T_{f}$ and interval sizes the negativity therefore decreases as the backreaction increases toward unity, i.e. the magnitude of this negative correction is proportional to $(1-\rho)$ and vanishes as $\rho\to 1$. Physically, this indicates that adding heavy-flavour backreaction tends to enhance thermal/classical correlations which compete with and reduce the purely quantum correlations measured by negativity in the low-temperature regime.
  • This behavior of negativity at low $T_{f}$ is consistent with the HEE analysis where backreaction-generated terms modify the finite part of the entropy and tend to push the system toward thermal-dominated behaviour; negativity—being a quantifier of distillable entanglement-like correlations—responds by decreasing when thermal contributions grow.
@@
The low-temperature adjacent-interval result (5.4) demonstrates that, in the near-vacuum regime, heavy-quark backreaction produces a leading thermal correction $\propto(1-\rho)T_{f}^{d}$ which lowers the HEN. This is consistent with the qualitative expectation that additional matter content increases thermal correlations and thereby diminishes quantum entanglement as measured by negativity.
@@ 5.2 High temperature limit
[…]
@@
$\mathcal{E}\;=\;\frac{3\,L^{d-2}R^{d-1}}{8G_{N}}\,\biggl(\frac{4\pi T_{f}}{d}\biggr)^{\!d-2}\,\Big[\,\widetilde{S}_{1}(\rho,d)+\widetilde{S}_{2}(\rho,d)-\widetilde{S}_{12}(\rho,d)\,\Big]\,$, (5.5)
@@
where the dimensionless functions $\widetilde{S}_{1},\widetilde{S}_{2},\widetilde{S}_{12}$ are defined in eq.˜3.14 and encode the backreaction dependence; numerically (and as established in [Chakrabortty:2020ptb]) each $\widetilde{S}(\rho,d)$ is a monotonically increasing function of $\rho$. From the above result, it is evident that
  • The combination of $\widetilde{S}$’s in square brackets is finite and captures all nontrivial dependence on the backreaction parameter $\rho$ and the spacetime dimension $d$. Divergent pieces present in the separate area integrals cancel in the prescribed linear combination, yielding the well-defined finite coefficient multiplying the thermal scaling.
  • Since each $\widetilde{S}(\rho,d)$ grows with $\rho$ as noticed in [Chakrabortty:2020ptb], the combination in (5.5) under squre bracket typically increases with $\rho$. Therefore, at high effective temperature the net effect of turning on quark backreaction is an increase of the finite part of $\mathcal{E}$.
  • Physically, this enhancement can be understood as a modification of the near-horizon geometry by the backreacting matter: the geometric measure that feeds into the negativity receives larger near-horizon contributions when $\rho$ is increased, which shows up as larger values of the $\widetilde{S}$ coefficients.
At high effective temperature the adjacent-interval HEN becomes an extensive thermal-like quantity and is enhanced by heavy-quark backreaction through the monotonic $\rho$-dependence of the coefficients $\widetilde{S}$.
[…]
6 HEN for bipartite subsystems
Having analyzed the adjacent subsystems configuration, we now extend our study to the bipartite case, where the subsystem $A$ and its complement $A^{c}$ together form the entire boundary theory. Unlike the previous setups, this configuration captures the global entanglement structure of the plasma and provides a more direct comparison between quantum and thermal correlations. In particular, the bipartite negativity is expected to be sensitive to the thermal entropy density at high temperature, while at low temperature it should predominantly reflect the underlying vacuum-like entanglement structure. Furthermore, the presence of backreaction introduces nontrivial corrections that allow us to probe how heavy quark density modifies the global entanglement content of the system.
@@ 6 HEN for bipartite subsystems
To compute the bipartite HEN, we divide the $A^{c}$ in two parts denoted by $B_{1}$ and $B_{2}$ on the either side of $A$. We can define the regions as,
$B_{1}\in[-\frac{l}{2},-L],\qquad A\in[-\frac{l_{1}}{2},\frac{l_{1}}{2}],\qquad B_{2}\in[\frac{l}{2},L].$ (6.1)
The HEN for this case is given by [Chaturvedi:2016rft]
$\mathcal{E}=\lim_{B\to A^{c}}\biggl[\frac{3}{16G_{N}}(2\mathcal{A}_{A}+\mathcal{A}_{B_{1}}+\mathcal{A}_{B_{2}}-\mathcal{A}_{A\cup B_{1}}-\mathcal{A}_{A\cup B_{2}})\biggr]$ (6.2)
Where $B=B_{1}\cup B_{2}$, we can choose $B_{1}=B_{2}$ then,
$\mathcal{E}=\lim_{B\to A^{c}}\frac{3}{8G_{N}}(\mathcal{A}_{A_{1}}+\mathcal{A}_{B_{1}}-\mathcal{A}_{A\cup B_{1}})$ (6.3)
where $B\to A^{c}$ is the bipartite limit and can be imposed as $L\to\infty$.
@@ 6.1 Low temperature analysis
Starting from the general prescription (6.3) and using the low-temperature expansions for the area functionals, one obtains for the bipartite configuration as
$\begin{split}\mathcal{E}&=\lim_{B\to A^{c}}\frac{3}{8G_{N}}\Bigg\{A_{0}+R^{d-1}S_{0}+R^{d-1}\rho\,S_{0}S_{2}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}\Big[\Big(\tfrac{L}{l}\Big)^{d-2}+\Big(\tfrac{L}{L-\tfrac{l}{2}}\Big)^{d-2}-\Big(\tfrac{L}{L+\tfrac{l}{2}}\Big)^{d-2}\Big]\\ &\qquad\qquad+R^{d-1}S_{0}S_{1}(1-\rho)\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d}\Big[\Big(\tfrac{L}{l}\Big)^{d-2}l^{d}+\Big(\tfrac{L}{L-\tfrac{l}{2}}\Big)^{d-2}\big(L-\tfrac{l}{2}\big)^{d}-\Big(\tfrac{L}{L+\tfrac{l}{2}}\Big)^{d-2}\big(L+\tfrac{l}{2}\big)^{d}\Big]\\ &\qquad\qquad+\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg\}\,.\end{split}$ (6.4)
Taking the large transverse-volume limit $L\to\infty$ (with $l$ held fixed and $L\gg l$) simplifies the geometric prefactors and yields the dominant large–$L$ form
@@
$\begin{split}\mathcal{E}=\frac{3}{8G_{N}}\Bigg\{A_{0}+R^{d-1}S_{0}+R^{d-1}S_{0}L^{d-2}\Big[\,l^{2-d}\;+\;\rho\,S_{2}\,l^{2-d}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}&\;+\;(1-\rho)\,S_{1}\,l^{4-d}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d}\;\Big]\\ &+\;\mathcal{O}\big((T_{f}l)^{2d-2}\big)\Bigg\},\end{split}$ (6.5)
where $A_{0}$ is the UV-divergent area-law piece and $S_{0},S_{1},S_{2}$ are $d$-dependent constants defined previously. One can draw the following conclusions from (6.5),
  • Within the finite part the leading geometric scaling is $l^{2-d}$ (the vacuum-like short-distance scaling for a strip), with two distinct backreaction corrections controlled by $\rho$.
@@
  • There are two $\rho$-dependent corrections with different parametric dependence: first is A $\rho$–linear correction. This term enters with the same geometric scaling $l^{2-d}$ as the leading finite term and therefore represents a multiplicative modification of the vacuum-like contribution. The second one is $(1-\rho)$. This carries an extra factor of $l^{2}$ relative to the first correction and is parametrically higher order in $T_{f}$.
@@
  • For the low-temperature expansion and for sufficiently small strip width $l$, the $\rho$–linear term (scaling as $T_{f}^{\,d-1}\,l^{2-d}$) typically dominates over the $(1-\rho)$ term.
This mixed behavior contrasts with the adjacent-interval low-$T_{f}$ result (where a $(1-\rho)T_{f}^{d}$ correction suppressed negativity); the bipartite configuration therefore exhibits a subtler interplay between competing $\rho$-dependent effects. The origin of this difference is geometric: the bipartite construction samples area contributions from large RT surfaces (via the $L^{d-2}$ factor), producing the additional $\rho S_{2}T_{f}^{d-1}$ multiplicative modification which is absent (or subleading) in the adjacent formulae.
6.2 High Temerature analysis
[…]

### v2

@@ 5 HEN for Adjacent Subsystems
$x\equiv x^{1}\in\biggl[-\frac{l_{j}}{2},\frac{l_{j}}{2}\biggr],\quad x^{i}\in\biggl[-\frac{L}{2},\frac{L}{2}\biggr];\quad j=1,2,$ (5.1)
where i=2,3,,d1i=2,3,...,d-1 and $L\to\infty$. The holographic prescription then requires an appropriate combination of the entanglement entropies of the individual regions and their union [40, 37]. This construction effectively isolates the quantum part of correlations by eliminating purely thermal contributions. In what follows, we will evaluate the holographic negativity for the adjacent strip configuration in the backreacted AdS black hole background and examine its dependence on quark density $\rho$, with particular focus on the low and high temperature limits, in order to contrast its behavior with that of the entanglement entropy discussed previously. Following the proposal given in [37], the HEN for the mixed state configuration of two adjacent subsystems is given by
$\mathcal{E}=\frac{3}{16G_{N}^{d+1}}(\mathcal{A}_{1}+\mathcal{A}_{2}-\mathcal{A}_{3}),$ (5.2)
Here, $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}=\mathcal{A}_{1}+\mathcal{A}_{2}$ represent the areas of the corresponding RT surfaces, obtained using the prescription reviewed in section˜3. To express the entanglement negativity in terms of boundary parameters, one needs to determine the turning point $z_{t}$ as a function of the strip width $l$ by solving (3.6). However, as discussed in section˜3, this relation cannot be solved exactly for arbitrary $z_{t}$ and admits analytic solutions only in the low and high temperature regimes. Therefore, we evaluate the HEN in these two limiting cases using the corresponding results for the HEE.
@@ 5.1 Low temperature limit
The low effective temperature regime corresponds to the condition $z_{t}\ll z_{H}$. In terms of the boundary parameter, this limit translates to $T_{f}l_{j}\ll 1$. By solving eq.˜3.6 perturbatively for $z_{t}$ in powers of $l$, one can obtain the corresponding expressions for the area functions $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ in this regime, as outlined in [14]. The resulting expressions for the area can be written from eq.˜3.9 is given by
@@
$\begin{split}\mathcal{A}_{j}=\frac{2}{d-2}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}&+R^{d-1}\left(\frac{L}{l_{j}}\right)^{d-2}S_{0}\Biggl[1+\rho S_{2}\left(\frac{4\pi T_{f}l_{j}}{d}\right)^{d-1}\\ &+(1-\rho)S_{1}\left(\frac{4\pi T_{f}l_{j}}{d}\right)^{d}+\mathcal{O}\left((T_{f}l_{j})^{2d-2}\right)\Biggr].\end{split}$ (5.3)
@@
Here, the index $j=1,2,3$ labels the subsystems $A_{1}$, $A_{2}$ and $A_{3}$ corresponding to their respective strip widths. In eq.˜5.3, the first term with $\epsilon$ denotes the divergent contribution while $S_{0}$, $S_{1}$, and $S_{2}$ are constants depending only on the spacetime dimension $d$ and are given in eq.˜3.10. Now using (5.2) together with the low temperature expansion in eq.˜5.3 for the area functionals, the HEN for two adjacent strip subsystems in the deformed AdS black hole background takes the following form
@@
$\begin{split}\mathcal{E}&=\frac{3}{8G_{N}^{d+1}(d-2)}\frac{L^{d-2}R^{d-1}}{\epsilon^{d-2}}+\frac{3R^{d-1}L^{d-2}}{16G_{N}^{d+1}}S_{0}\Biggl[l_{1}^{2-d}+l_{2}^{2-d}-(l_{1}+l_{2})^{2-d}\\ &\qquad\qquad\qquad\qquad\qquad\qquad-\;2S_{1}(1-\rho)\,l_{1}l_{2}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{\!d}+\mathcal{O}\!\big((T_{f}l_{j})^{2d-2}\big)\Biggr].\end{split}$ (5.4)
The following observations can be made from the above expression in eq.˜5.4 as follows:
@@
  • The first term is the standard ultraviolet divergence arising from short distance correlations near the entangling surface. The leading finite contribution (independent of $T_{f}$) is the familiar vacuum like short distance scaling for adjacent strip geometries and encodes the purely geometric contribution to negativity.
@@
  • The first nontrivial thermal/backreaction correction appears at order $T_{f}^{d}$ and is multiplied by the factor $(1-\rho)$. In low temperature regime $T_{f}^{\,d}$ term is the leading thermal correction and higher order terms scale as $(T_{f}l_{j})^{2d-2}$ which are subleading for sufficiently small $T_{f}l_{j}$.
  • The correction term proportional to $(1-\rho)T_{f}^{\,d}$ increases negativity. For fixed $T_{f}$ and subsystem sizes, the HEN therefore increases as the backreaction increases. Physically, this indicates that adding heavy flavor backreaction tends to enhance the correlations due to the addition of extra degrees of freedoms coming from the addition of backreaction.
@@
The low temperature adjacent subsystems result in eq.˜5.4 demonstrates that in the near-vacuum regime, heavy-quark backreaction produces a leading thermal correction that raise the HEN. This is consistent with the qualitative expectation that additional matter content increases correlations and thereby enhance quantum entanglement as measured by negativity. We also observe that for $\rho=0$, the expression in eq.˜5.4 simplifies to the case of an undeformed geometry and matches with the corresponding HEN for adjacent subsystems at low temperature limit [37].
@@ 5.2 High temperature limit
In the regime of high effective temperature i.e when $T_{f}l_{j}\gg 1$, the HEN for the mixed state configuration of adjacent subsystems can be obtained using eq.˜3.13 and eq.˜5.2 as follows
@@
$\mathcal{E}=\;\frac{3\,L^{d-2}R^{d-1}}{8G_{N}^{d+1}}\left[\frac{1}{(d-2)\epsilon^{d-2}}+\biggl(\frac{4\pi T_{f}}{d}\biggr)^{\!d-2}\,\widetilde{S}(\rho,d)\right],$ (5.5)
@@
where the dimensionless functions $\widetilde{S}$ is defined in eq.˜3.14 and encode the backreaction dependence. In [14] it has been shown that $\widetilde{S}(\rho,d)$ is a monotonically increasing function of $\rho$. From eq.˜5.5, we observe the following:
  • Unlike the HEE, the HEN is free from any volume dependent contribution. The terms proportional to the subsystem volume which encode classical and thermal correlations cancel exactly as expected. Since entanglement negativity quantifies purely quantum correlations in mixed states, it does not exhibit volume scaling and follow the area law behavior.
  • The quantity $\widetilde{S}$ is finite and encodes the entire nontrivial dependence on the backreaction parameter $\rho$ and the spacetime dimension $d$. As $\widetilde{S}$ increases monotonically with $\rho$, the holographic entanglement negativity correspondingly exhibits an enhancement with increasing backreaction.
  • Physically, this enhancement originates from the modification of the near horizon geometry due to matter backreaction. As the backreaction parameter $\rho$ increases, the geometric contribution entering the entanglement negativity acquires larger near horizon corrections which is reflected in the growth of $\widetilde{S}$.
In the high temperature regime, the HEN retains a single ultraviolet divergence matching the expected short distance behavior while remaining free from any volume law contribution due to the exact cancellation of volume dependent terms. The entire nontrivial dependence on the backreaction parameter $\rho$ and the spacetime dimension $d$ is encoded in the finite quantity $\widetilde{S}$ which increases monotonically with $\rho$. This enhancement reflects the growing influence of matter backreaction on the near horizon geometry leading to larger geometric contributions to the entanglement negativity. In the next section we will analyze the HEN for bipartite configuration.
6 HEN for Bipartite System
Having analyzed the adjacent subsystems configuration, we now turn to the bipartite configuration, where the subsystem $A$ and its complement $A^{c}$ together form the entire boundary theory. In this setup, holographic entanglement negativity provides a sharp probe of mixed state quantum correlations allowing us to distinguish genuine quantum entanglement from classical and thermal correlations that dominate at high temperature.
@@ 6 HEN for Bipartite System
To compute the HEN for bipartite system, we divide the $A^{c}$ in two parts denoted by $B_{1}$ and $B_{2}$ on the either side of $A$. We can define the subsystems as
$B_{1}\in[-\frac{l}{2},-L_{1}],\qquad A\in[-\frac{l}{2},\frac{l}{2}],\qquad B_{2}\in[\frac{l}{2},L_{2}].$ (6.1)
The HEN for this mixed state configuration is then given by [19]
$\mathcal{E}=\lim_{B\to A^{c}}\biggl[\frac{3}{16G_{N}^{d+1}}(2\mathcal{A}_{A}+\mathcal{A}_{B_{1}}+\mathcal{A}_{B_{2}}-\mathcal{A}_{A\cup B_{1}}-\mathcal{A}_{A\cup B_{2}})\biggr],$ (6.2)
where $B=B_{1}\cup B_{2}$. We can choose $B_{1}=B_{2}$ which leads to the following
$\mathcal{E}=\lim_{B\to A^{c}}\frac{3}{8G_{N}^{d+1}}(\mathcal{A}_{A}+\mathcal{A}_{B_{1}}-\mathcal{A}_{A\cup B_{1}}),$ (6.3)
where $B\to A^{c}$ is the bipartite limit and can be imposed as $L_{1}\to\infty$.
@@ 6.1 Low temperature analysis
In this section, we compute the HEN for a bipartite configuration in the low effective temperature limit defined by $T_{f}l\ll 1$. In the bulk, this regime implies that the turning point of the minimal surface for subsystem $A$ is located far from the black hole horizon. Conversely, for the subsystems $B_{1}$ and $A\cup B_{1}$ with lengths $(L_{1}-l/2)$ and $(L_{1}+l/2)$, the limit $B\to A^{c}$ $(L_{1}\to\infty)$ ensures that their respective minimal surfaces $\mathcal{A}_{B_{1}}$ and $\mathcal{A}_{A\cup B_{1}}$ probe the deep bulk and approach the horizon. By substituting the low temperature area $\mathcal{A}_{A}$ from eq.˜3.11 and the high temperature limits for $\mathcal{A}_{B_{1}}$ and $\mathcal{A}_{A\cup B_{1}}$ from eq.˜3.13 into eq.˜6.3, we obtain the HEN for bipartite configuration as follows
@@
$\begin{split}\mathcal{E}=\frac{3R^{d-1}}{8G_{N}^{d+1}}\Bigg[\frac{2}{(d-2)}\left(\frac{L}{\epsilon}\right)^{d-2}+S_{0}\left(\frac{L}{l}\right)^{d-2}+L^{d-2}\rho\,S_{0}S_{2}l\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}\\ +L^{d-2}S_{0}S_{1}(1-\rho)l^{2}\Big(\frac{4\pi T_{f}}{d}\Big)^{d}-V\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}\Bigg],\end{split}$ (6.4)
where $V=lL^{d-2}$ is the volume of the subsystem $A$ in a $d$ dimensional dual field theory. The above expression for the HEN in the low effective temperature regime can be written in the following form
$\mathcal{E}=\frac{3}{2}\left(S_{A}-S_{A}^{th}\right),$ (6.5)
where $S_{A}$ is the entanglement entropy for the subsystem $A$ given in eq.˜3.11 and $S_{A}^{th}=\frac{R^{d-1}}{4G_{N}^{d+1}}\Big(\tfrac{4\pi T_{f}}{d}\Big)^{d-1}V$ is the thermal entropy associated with the subsystem $A$. The above relation in eq.˜6.5 demonstrates that at leading order, the HEN is obtained by subtracting the volume dependent thermal entropy from the HEE for the finite temperature mixed state of a single subsystem. This result aligns with quantum information theoretic expectations where entanglement negativity serves as an upper bound on distillable entanglement by isolating the purely quantum correlations in a mixed state. This kind of relation has also been established in the literature at various scenarios [19, 65, 45]. One can draw the following conclusions from the above result:
  • The first term represents the UV divergent contribution similar to HEE. Within the remaining finite part, the dominant geometric scaling is governed by $l$ which recovers the characteristic short distance behavior of a strip in a pure AdS geometry.
@@
  • Unlike the adjacent case there are two $\rho$-dependent corrections with different parametric dependence first is a $\rho$-linear correction that increses HEN with backreaction. This term enters with the same geometric scaling $l$ as the leading finite term. The second one is $(1-\rho)$ and carries a factor of $l^{2}$ relative to the first correction and is parametrically higher order in $T_{f}$. This term reduces the negativity as backreaction increases.
@@
  • For the low temperature expansion and for sufficiently small strip width $l$, the $\rho$ linear term typically dominates over the $(1-\rho)$ term.
Overall, the backreaction effects enter through two distinct $\rho$-dependent corrections: a linear term that multiplicatively modifies the vacuum contribution and a subleading $(1-\rho)$ term that is parametrically suppressed for small strip widths and low temperatures. In this regime, the $\rho$-linear correction provides the dominant backreaction effect on the HEN.
6.2 High temperature analysis
In the limit of high effective temperature $(T_{f}l\gg 1)$, we combine the bipartite prescription (6.3) with the area expansions from eq.˜3.13 to arrive at the following expression for the HEN as